### v2

# Speculative Diffusion Decoding: Accelerating Language Generation through Diffusion

Abstract
Speculative decoding has emerged as a widely adopted method to accelerate large language model inference without sacrificing the quality of the model outputs. While this technique has facilitated notable speed improvements by enabling parallel sequence verification, its efficiency remains inherently limited by the reliance on incremental token generation in existing draft models. To overcome this limitation, this paper proposes an adaptation of speculative decoding which uses discrete diffusion models to generate draft sequences. This allows parallelization of both the drafting and verification steps, providing significant speed-ups to the inference process. Our proposed approach, Speculative Diffusion Decoding (SpecDiff), is validated on standard language generation benchmarks and empirically demonstrated to provide a up to 8.7x speed-up over standard generation processes and up to 2.5x speed-up over existing speculative decoding approaches.
Keywords Parallel Decoding  $\cdot$ Large Language Models  $\cdot$ Discrete Diffusion Models
1 Introduction
As autoregressive language modeling with transformers (Vaswani et al., 2017) is scaled to larger compute levels, performance improves and new capabilities emerge (Kaplan et al., 2020; Brown et al., 2020). Indeed, scaling large language models (LLMs) makes them helpful to broad audiences for code generation, question answering, summarization, and other use cases (Achiam et al., 2023; Gemini Team, 2023; Llama Team, 2024), motivating the deployment and public release of increasingly large models. However, running LLMs in inference mode for millions of users produces burdensome electricity, time, and monetary demands. Many methods exist to mitigate these costs – including sparsity, quantization, and distillation – but they often introduce new problems (e.g., their application can degrade the performance of the model) (Hong et al., 2024).
[…]
Notably, since both the drafting model’s speed and its quality relative to the target model are critical to the success of speculative decoding, simultaneous improvements in each of these areas are necessary to ensure speculative decoding’s relevance to future, more capable target models. For instance, a small GPT-2 (Radford et al., 2019) drafting model could produce drafts that are often rejected by GPT-4 (Achiam et al., 2023), and simply scaling the drafting model to address its weaker generations risks diminishing the speed advantage necessary to speculative decoding’s success.
To address this challenge, this paper proposes a method that leverages the recently introduced discrete diffusion models (Lou et al., 2024). These models offer a smooth tradeoff between the compute cost of generation and the quality of generation (via the number of reverse diffusion steps). Moreover, while they have historically struggled relative to traditional language models, recent diffusion models have been shown to require 32×32$\times$32 × fewer function evaluations than autoregressive models to produce text with comparable perplexity (Lou et al., 2024). Additionally, future advances in diffusion model generation quality are highly aligned with their ability to perform strongly as speculative drafters: as drafted tokens are accepted by the target model at a higher rate, a larger number of proposed drafted tokens becomes optimal from an efficiency/speed point of view, and (unlike sequential drafters) diffusion models can easily accommodate generation of more tokens since they are able to generate entire sequences in one step.
Contributions. More specifically, this paper makes the following contributions: (1) It introduces a novel integration of generative diffusion language models with speculative decoding, schematically illustrated in Figure 1 (right). (2) It empirically demonstrate the hybrid model’s ability to significantly accelerate inference times while maintaining the same high-quality outputs of the original, target large language model. (3) The proposed method ensures that all generations from the diffusion language model, which are empirically shown to produce significantly worse outputs than current state-of-the-art autoregressive models (Lou et al., 2024; Austin et al., 2021; Gloeckle et al., 2024), align with the outputs generated by larger, more computationally demanding models. (4) Finally, the paper sets a new benchmark for speed in language completion tasks on the CNN/DM and OpenWebText datasets.
Figure 1: Schematic illustration of classical speculative decoding (left) and speculative diffusion decoding (right).
2 Related Work
While autoregressive language models provide state-of-the-art performance on language generation tasks, the incremental decoding used by these architectures results in significant overhead at inference time (Miao et al., 2023a). This is largely a result of the inability to parallelize the sequential process of generating tokens in the output sequence as each token generation is dependent upon the preceding tokens in the sequence; consequentially, scaling the compute associated with the inference cannot directly reduce this overhead when using standard decoding schemes. In recent literature studying how to accelerate large language model generation, two primary approaches have been explored: (1) advanced decoding implementations that better parallelize token generation and (2) non-autoregressive language models allowing full sequences to be generated simultaneously.
Speculative decoding. accelerates autoregressive generation by leveraging a smaller autoregressive models of the same architecture (the drafter model) to predict candidate sequences for the original model (the target model) to verify (Leviathan et al., 2023; Chen et al., 2023). Notably, the earliest literature on speculative diffusion adapted a non-autoregressive model to act as the drafter model (Xia et al., 2023), utilizing a masked language model with a bidirectional decoder (Ghazvininejad et al., 2019). However, the integration of non-autoregressive draft models has not received much attention due to the difficulty introduced by the necessary additional training in existing approaches and the modest speed-ups that were previously reported using these methods (less than 2x speed-up over vanilla decoding schemes).
Recent advancements in speculative decoding have focused on overcoming memory-related constraints, with improvements achieved through various approaches: drafting directly with the target model (Cai et al., 2024; Zhang et al., 2024), enhancing draft algorithms (Sun et al., 2024), and introducing additional parallelization techniques that incorporate branching to refine the drafting process (Fu et al., 2024; Miao et al., 2023b; Svirschevski et al., 2024).
Non-autoregressive language models. Models which stray from the autoregressive paradigm have been shown to speed-up generation by generating blocks or even entire sequences simultaenously. Gloeckle et al. (2024) propose a method of adapting traditional autoregressive models to sample blocks of tokens, improving inference time over similarly scaled models. In a similar vein, diffusion language models have been recognized for their efficiency in generating extended token sequences concurrently, offering even greater speed enhancements. These models recast language generation as a diffusion process either across the embedding space (Austin et al., 2021) or, more recently, through the probability distributions of generated tokens (Lou et al., 2024). Current, state-of-the-art models report up to a 32x speed-up over similarly sized GPT-2 models Lou et al. (2024). However, despite the fact that these models have been shown to dramatically accelerate the inference time for language generation, diffusion models typically perform less effectively than state-of-the-art autoregressive models in terms of standard language metrics, often exhibiting significantly higher perplexity scores. In the following section, we will demonstrate, for the first time, how the speed of these models can be leveraged without being subject to this critical limitation.
3 Preliminaries and Settings
We start by formalizing the settings and goals. For open-ended language generation, we focus on the task of token generation, where given a sequence of tokens $x_{1},x_{2},\ldots,x_{i}$, the goal is to generate the next $n$ tokens $x_{i+1},\ldots,x_{i+n}$ from the conditional distributions $p(x_{i+1}|x_{1},x_{2},\ldots,x_{i}),\ldots,p(x_{i+n}|x_{1},x_{2},\ldots,x_{i+n-1})$ or more succinctly $p_{i+1},\ldots,p_{i+n}$.
Speculative decoding. Speculative decoding leverages two LLMs, $M_{p}$ and $M_{q}$, to parallelize token generation:
  • $M_{p}$ is the original, target, model whose output probability distributions for the tokens are $p_{i+1},\ldots,p_{i+n}$.
  • $M_{q}$ is a smaller and more efficient drafter model, used to generate approximations of the distribution of $M_{p}$ as $q_{i+1},\ldots,q_{i+n}$.
This process follows a draft-then-verify approach (Stern et al., 2018), where $M_{q}$ efficiently computes a candidate sequence of tokens, which $M_{p}$ then verifies in parallel.
During each speculative decoding iteration, $M_{q}$ generates a subset of the total $n$ tokens that are required for the generation task. The size of this subset is denoted as $\gamma$. As shown in Figure 1 (left), the tokens $x_{i+1},\ldots,x_{i+\gamma}$ sampled from $M_{q}$ are then used by $M_{p}$ to generate the corresponding probability distributions $p_{i+1},\ldots,p_{i+\gamma}$. The distributions $q_{i+1},\ldots,q_{i+\gamma}$ from $M_{q}$ are stored for evaluating acceptance in subsequent steps. Critically, the target model’s inference over $p_{i+1},\ldots,p_{i+\gamma}$ can now be run in parallel as the model has access to tokens $x_{i+1},\ldots,x_{i+\gamma}$, alleviating the sequential dependency for generation with $M_{p}$.
To ensure high-quality outputs despite potential discrepancies between $M_{p}$ and $M_{q}$, tokens are subjected to an acceptance criterion. For each token $x_{i+1},\ldots,x_{i+\gamma}$, if $q(x)\leq p(x)$, the token is accepted. If $q(x)>p(x)$, the token is rejected with a probability of $1-\frac{p(x)}{q(x)}$. This criterion is applied sequentially from left to right; rejection of any token results in the discard of all subsequent tokens. Hence, the token acceptance is maximized when the output distributions of $M_{q}$ and $M_{p}$ are closely aligned. Previous literature quantifies the likelihood of token acceptance, denoted $\alpha$, and theoretically demonstrate that $\alpha=1-\mathbb{E}(D_{LK}(p,q))$ where $D_{LK}$ represents the divergence between the distributions Leviathan et al. (2023). This has led to the prevalent use of drafters taken from the same series as the target models, a paradigm that we challenge in this paper.
4 Speculative Diffusion Models
Overview.
Speculative decoding has provided state-of-the-art results for improving language generation inference time but requires meticulous tuning of the associated hyperparameters to achieve optimal results. Particularly $\gamma$, the sequence length generated by the drafter model, needs to be appropriately calibrated not only to maximize potential speed-up but to even outperform standard autoregressive decoding. This is an important consideration when using current autoregressive draft models, provided that the inference time to generate $M_{q}(x)$, the draft logits, is directly scaled by the size of $\gamma$. Increasing this value too high reduces the number of operations that are conducted in parallel, potentially leading to speculative decoding increasing inference time, while reducing this value too low results in speculative decoding “missing out” on token generations that could have been handled by the draft model.
Leviathan et al. (2023) has conducted theoretical analysis on how to best optimize the value of $\gamma$, however, it has been contingent upon accurately estimating the percentage of tokens in a the sequence that will be accepted by the target model. By their own acknowledgment, it would be necessary to predict this value for each draft and numerically solve for the optimal value of $\gamma$ to fully realize the potential speed-up of speculative decoding. Thus, a significant portion of the residual suboptimality in current implementations can be attributed directly to the sensitivity of this hyperparameter.
Diffusion language models are juxtaposed to conventional language models in that they do not sample token sequences in a consecutive manner, rather generating entire sequences in parallel. This has resulted in significant speed-up over similarly sized autoregressive models when generating extended sequences (Lou et al., 2024). This can particularly be observed in longer sequence generations as scaling the draft length $\gamma$ results in minimal overhead due to the ability to directly parallelize token generation.
In the following section, we empirically demonstrate how these attributes of diffusion language models – combined with the intrinsic speed gains from recent advancements in diffusion techniques – position these models as exceptionally promising candidates for drafting within a speculative decoding framework.
4.1 SpecDiff: Formulation
For language modeling, discrete diffusion models enable the diffusion process to generate sequences that fit within a combinatorial output space. Unlike continuous diffusion models, traditional score-matching techniques cannot be directly applied to learn discrete diffusion models. Instead, various surrogate losses have been proposed for training. Notably, Lou et al. employ a score entropy loss that models the ratio between the probability mass vectors of the noisy distribution, $p(x)$, and an interim distribution closer to the training data, $p(y)$:
$\mathbb{E}_{x\sim p}\left[\sum_{y\neq x}w_{xy}\left(s_{\theta}(x)_{y}-\frac{p(y)}{p(x)}\log s_{\theta}(x)_{y}+K\left(\frac{p(y)}{p(x)}\right)\right)\right]$ (1)
where $s_{\theta}(x)_{y}$ is the diffusion model, $w$ is a matrix of non-negative weighting values, and $K(a)=a(\log a-1)$ is a normalizing function. This loss is directly used for pretraining and finetuning of our draft model. As this process learns to denoise over the probability mass vectors, the output of the draft model is a matrix of $\mathbb{R}^{n\times m}$ where $n=\gamma$ and $m$ is the size of the vocabulary. The candidate sequence can
$\triangleright$ Take $T$ diffusion steps to generate the draft.
$q_{i+1,\ldots,i+\gamma}^{T}\sim\mathcal{N}(\mathbf{0},\sigma_{T}\mathbf{I})$
for $t=T$ to 1 do
       $q_{i+1,\ldots,i+\gamma}^{t-1}(x)\leftarrow M_{q}([x_{0},\ldots,x_{i}]+[q_{i+1,\ldots,i+\gamma}^{t}(x)],t)$
$x_{i+1,...,i+\gamma}\sim{q}^{0}$
$\triangleright$ Run $M_{p}$ in parallel.
$p_{i}(x),\ldots,p_{i+\gamma+1}(x)\leftarrow$   $M_{p}(x_{0},\ldots,x_{i}),\ldots,M_{p}(x_{0},\ldots,x_{i+\gamma})$
$\triangleright$ Determine the number of accepted guesses $n$.
$r_{i}\sim U(0,1),\dots,r_{i+\gamma}\sim U(0,1)$
$n\leftarrow\min(\{j-1\mid i\leq j\leq i+\gamma,r_{j}>\frac{p_{j}(x)}{q_{j}(x)}\}\cup\{\gamma\})$
$\triangleright$ Adjust the distribution from $M_{p}$ if needed.
$p^{\prime}(x)\leftarrow p_{n+1}(x)$
if $n<i+\gamma$ then
       $p^{\prime}(x)\leftarrow norm(max(0,p_{n+1}(x)-q_{n+1}(x)))$
$\triangleright$ Return one token from $M_{p}$, $n$ tokens from $M_{q}$.
$t\sim p^{\prime}(x)$
return $x_{1},\ldots,x_{n},t$
Algorithm 1 SpecDiff Decoding
then be generated using standard decoding methods over these probability mass vectors, the logits of which are stored to be used when determining whether to accept each draft token.
Now, the draft logits produced by the output matrix of the discrete diffusion drafter directly substitue the autoregressive drafter used to generate $M_{q}([x_{0},\ldots,x_{i}]+[x_{i+1},\ldots,x_{i+\gamma}])$. This substitutes the draft step taken by Leviathan et al. and Chen et al. This is the primary difference between SpecDiff and standard speculative decoding approaches, and subsequent steps of verifying this draft with the target model follow the previously proposed decoding algorithm. We provide a complete overview of the SpecDiff decoding in Algorithm 1 (adapted from (Leviathan et al., 2023)).
We highlight that while in standard speculative diffusion the number of evaluations by the drafter model is dictated by the value of $\gamma$ (used in the first loop for Algorithm 1), in our implementation it is dictated by the number of diffusion steps, $T$. This allows SpecDiff to scale $\gamma$ to much higher values as discussed further in Section 5.2. Instead, the value of $T$ is selected to optimize the trade-off between draft quality and computational overhead. While analysis by Lou et al. shows that lower values of $T$ lead to higher perplexity in the generated sequence, this only impacts SpecDiff with respect to its effect on the percentage of tokens from the draft which are accepted (Figure 2).
5 Experiments
To empirically evaluate the improvements provided by utilizing SpecDiff, analysis is provided on text summarization and more general text generation, leveraging benchmarks that are common to existing literature. All evaluation is conducted on two NVIDIA A100 series GPUs (80GB) utilizing CUDA 11.8. Additionally, FlashAttention Dao et al. (2022) is used to optimize the performance in all experiments.
5.1 Experimental Setup
Settings.
Evaluation is conducted on two standard natural language processing tasks: (1) text summarization using the CNN/DM dataset and (2) text generation finetuned on the OpenWebText dataset. In each setting the model is queried for 1024 tokens using a probabilistic decoding scheme (temperature = 1). For each experiment we evaluate the target models $M_{p}$ GPT-2 XL (1.5B) and GPT-NEO (2.7B) finetuned on the specified dataset with drafter models $M_{q}$ finetuned SEDD-Absorbing Small (90M) which is a comparable size to the baseline drafter GPT-2 (86M) Lou et al. (2024). The target model architectures are selected based on the criteria that they are the largest models that utilize a common tokenizer to pretrained SEDD weights publicly available.
Evaluation metrics.
Our method is assessed empirically by walltime speed-up and accepted tokens per draft, the product of $\alpha$, the ratio of accepted tokens, and $\gamma$. The latter metric is particularly relevant in this setting, as opposed to optimizing $\alpha$, given we empirically demonstrate the negligible overhead to extending $\gamma$ when using a diffusion-based drafter. The reported results are compared to recognized baselines of vanilla autoregressive decoding and standard speculative decoding implementations as proposed by Leviathan et al.; Chen et al., which our method most closely resembles.
$M_{p}$ $M_{q}$ $\gamma$ $\alpha$ Speed Up
CNN/DM Spec GPT-2 XL GPT-2 5 0.80 2.95x
GPT-NEO GPT-2 5 0.81 4.50x
Ours GPT-2 XL SEDD Small 40 0.89 7.59x
GPT NEO SEDD Small 40 0.91 8.73x
OpenWeb Spec GPT-2 XL GPT-2 5 0.76 3.16x
GPT-NEO GPT-2 5 0.77 3.38x
Ours GPT-2 XL SEDD Small 45 0.89 4.79x
GPT NEO SEDD Small 45 0.92 5.50x
Table 1: Evaluation of walltime speedup over autoregressive decoding using SpecDiff (ours) compared to standard speculative decoding (spec). The best result for each setting and target model is displayed in bold.
Table 2: Accepted tokens per draft comparing SpecDiff (ours) to standard speculative decoding (spec).
5.2 Results and Discussion
The following demonstrate the improvement provided utilizing a diffusion-based drafter model using the SpecDiff paradigm. Table 2 highlights the comparison between this approach and the use of an autoregressive drafter model. Across the tested settings and target model architectures, SpecDiff significantly outperforms standard speculative decoding methods, achieving speed-ups of up to 8.7x compared to the target models and increasing the efficiency of speculative decoding by more than 2.5x.
While previous implementations of speculative decoding rely on a common architecture between the drafter and target models Leviathan et al. (2023); Chen et al. (2023), using smaller versions of the same architecture to generate draft sequences, these experiments demonstrate a robustness to utilizing a completely different architecture for sequence drafting. While the acceptance rate $\alpha$ decreases when using SpecDiff, this is largely because of the increased $\gamma$, and thus the number of tokens accepted per draft sequence has still significantly increased. The much larger values of $\gamma$ used for SpecDiff should particularly be highlighted. This is a key discrepancy between diffusion language models, which generate entire sequences in parallel, and autoregressive models. Hence, there is minimal overhead to increasing the sequence length generated by the diffusion-based drafter, and $\gamma$ can be significantly increased without incurring significant cost. As a result, the number of tokens accepted during each generation increases (Table 2) despite the lower acceptance rate (reported as $\alpha$ in Table 2). This directly contributes to the improved performance of SpecDiff.
Figure 2: Evaluation of SpecDiff’s sensitivity to $\gamma$ and number of diffusion steps when optimizing speed (left) and accepted tokens per draft (right) as reported on the OpenWebText task using GPT-2 XL as the target model.
The hyperparameters used in the reported results have been optimized empirically. We highlight that while in standard speculative diffusion the performance is highly sensitive to $\gamma$, SpecDiff is robust to a range of values for $\gamma$ making it unnecessary to precisely tune this hyperparameter (in our experiments we found between 40 and 45 worked well). Rather, SpecDiff’s performance is much more sensitive to the number of diffusion steps selected. Similar to the role of $\gamma$ in an autoregressive model, the number of diffusion steps $T$ dictates the number of network evaluations during a single drafting step. As reported in the Figure 2, while increasing this hyperparameter arbitrarily results in higher values of $\alpha$, SpecDiff performs best when this is optimized to balance the objectives of maximizing the number of accepted tokens and minimizing the drafter’s overhead.
6 Future Work and Limitations
While this paper proposes a significant advance in the speculative decoding literature, SpecDiff motivates further work in this area as we hold the integration of diffusion language models with autoregressive models to be foundational. The current implementation of SpecDiff is limited to models which use the GPT-2 tokenizer, leveraging the pretrained SEDD models which have been trained with this, and adapting this to larger models will likely result in further speed improvements over standard speculative decoding. Furthermore, this paper has not fully realized improvements that could be made by hot-starting the drafter model with the logits of rejected tokens, as using partially generated information has already been shown to be effective when using diffusion models of different modalities Ruhe et al. (2024).
Additionally, our evaluation is limited in its comparison to standard speculative decoding implementations (Leviathan et al., 2023; Chen et al., 2023). More recent research on this topic has proposed methods which utilize additional parallelism to introduce branching in the drafting process, extending the length of accepted sequences by generating candidates sequences for the top-k tokens at any given point in the sequence (Fu et al., 2024; Miao et al., 2023b; Svirschevski et al., 2024). This work chooses not to compare to these methods as SpecDiff is complementary to such implementations, as the use of a diffusion-based drafter is in no way opposed to these approaches. In further studies we intend to extend SpecDiff to tree-based speculative decoding schemes to demonstrate its utility in the massively parallel settings explored by these works.
We also note that SpecDiff performs best on longer generation tasks. In Appendix A results on a shorter generation task, generations of less than 100 tokens, are provided using the LM1B dataset. While SpecDiff still outperforms standard speculative decoding on shorter generation tasks, the margin of improvement is reduced as it becomes impractical to scale $\gamma$ as dramatically. This is a byproduct of the dataset, where often the responses that the model is fine-tuned on are less than 30 tokens. Hence, adjusting $\gamma$ above this value provides no added benefit. Thus, although these are less practical settings, it is nevertheless important to noice that they would reduce the efficacy of SpecDiff by limiting draft model parallelization.
7 Conclusion
Motivated by the costly inference time of current large language models, this paper has proposed the novel integration of discrete diffusion models with autoregressive language models. The proposed method, Speculative Diffusion Decoding, alters existing speculative decoding schemes to integrate a non-autoregressive diffusion model as the draft model. As shown by the empirical evaluation on standard language generation benchmarks, the proposed method leverages the dramatic runtime advantages of recent work in discrete diffusion models for language generation while also maintaining the dramatically higher quality of an autoregressive target model. The reported results demonstrate the utility of this approach in effectively accelerating runtime, outperforming vanilla decoding by over 8x and speculative decoding methods by over 2.5x.
8 Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 and was supported by the LLNL-LDRD Program under Project No. 24-ERD-010. The work is also partially supported by NSF awards 2143706, 2232054, and 2242931. The views and conclusions provided in this paper reflect those of the authors only.
Appendix A Additional Results
To augment the longer sequence generation tasks that have been explored, additional experiments were conducted on the LM1B dataset. For this text generation task, blocks of 100 tokens were generated as outputs to a given prompt.
$M_{p}$ $M_{q}$ $\gamma$ $\alpha$ Speed Up
LM1B Spec GPT-2 XL GPT-2 5 0.69 1.70x
GPT-NEO GPT-2 5 0.81 2.47x
Ours GPT-2 XL SEDD Small 30 0.87 2.35x
GPT NEO SEDD Small 30 0.86 3.25x
Table 3: Evaluation of walltime speedup over autoregressive decoding using SpecDiff (ours) compared to standard speculative decoding (spec) on the LM1B dataset.
The results provide several interesting insights. First, we note that the optimal values of $\gamma$ are much lower for shorter generations. We hypothesize that this is directly due to the more succinct generation lengths. We also note that this could be impacted by the fine-tuning data, as the sequence lengths learned on for this dataset are generally much shorter than the other data settings explored. Second, it should be highlighted that in spite of these limitations SpecDiff provides a fair speed-up over the baseline, despite this being a generation task that does not exhibit the strengths of discrete diffusion models. It should be noted that in all experiments conducted on the SEDD models by Lou et al. generation lengths are explicitly set to 1024. We suspect this decision was intentional, as it best highlights the performance of these models, and acknowledge shorter generations to be a current limitation of discrete diffusion models.